There is a 50 meter long train. It magically goes at 60% the speed of light (0.6c). How long will it appear to be to an outside observer?


1
Expert's answer
2022-01-13T09:20:16-0500

Answer

Train length


L=50m


Moving v=0.6c


c is speed of light


So train will appear to be to an outside observer as

"L'=L\sqrt{1-\frac{v^2}{c^2}}"

"L'=50\sqrt{1-\frac{(0.6c)^2}{c^2}}"

"L'=50\sqrt{0.64}=40m"